John’s demand function for chocolate is Q=10-2p His price elasticity of demand for chocolate at a price p= -2/3 Find the value of p
Answer #1

Answer

elasticity of demand =(dQ/dP)*(P/Q)

dQ/dP=-2 ........ the first derivate of the demand function

using it

-2/3=(-2)*(P/(10-2P)) ......... using Q=10-2P

10*(2/3)- (2P)*(2/3)=2P

20/3 - 4P/3=2P

20-4P=6P

6P+4P=20

10P=20

P=2

the price is $2

---------